a bag 10 white, 15 red and 8 green balls. A single draw of 3 balls is made. what is the probability that a white, a red and green balls are drawn


1
Expert's answer
2021-12-13T16:29:15-0500

"P = \frac{C^{10}_1 \times C^{15}_1 \times C^8_1}{C^{33}_1} \\
= \frac{10 \times 15 \times 8}{5456} \\
= 0.2199"